OpenStudy (anonymous):

Using the identity Cos(A+B)=CosACosB+SinASinB Prove that Cos2A=1-2Sin^2A

OpenStudy (gorv):

put B=A cos2A=cosA.cosA+sinA.sinA

OpenStudy (anonymous):

Sorry top bit should be Cos(A+B)=CosACosB-SinASinB

OpenStudy (gorv):

sry then also do the same

OpenStudy (gorv):

(cosA)^2+sinA)^2=1-(sina)^2-(sina)^2

mathslover (mathslover):

\(\large {\cos 2A = \cos (A+A) = \cos A \cos A - \sin A \sin A \\ \implies \cos 2A = \cos^2 A - \sin^2 A \\ \implies \cos 2A = (1-\sin^2 A ) - \sin^2 A ... \quad \textbf{as : } \space \cos^2 A = 1 - \sin^2 A \\ \implies \cos 2A = 1 - 2\sin^2 A ... \boxed{\textbf{Proved}}} \)

OpenStudy (anonymous):

Isnt it\[\cos ^{2}A-1=1-\sin ^{2}A\] ?

mathslover (mathslover):

No, see : \(\large \cos^2 A + \sin^2 A= 1\) So, \(\large \cos^2 A = 1 - \sin^2 A\) \(\implies \large{\cos^2 A - 1 = -\sin^2 A}\)

OpenStudy (anonymous):

Ok I understand, my maths formula sheet was a bit confusing. Thanks for the help!

mathslover (mathslover):

You're welcome. Best of Luck and let me know if you have any more questions any time.